Two minibikes simultaneously left at the same time from certain country club and headed in opposite directions. At the end of an hour and 40 minutes they were 100 km apart. If one of the bikes traveled 5 kph faster that the other, find the speed of each bike?
Answer by mananth(16949)

mini bike I x mph
mini bike II x+ 5
..
They are going away from each other.

So add up their speed.

combined speed = x+ 1 x + 5
( 2 x + 5 )
Time = 1 2/ 3 hours
Distance = 100 miles
Distance = speed * time
( 2 x 5 )* 12/31/99 = 100
3.33 x + 8.33 = 100
3.33 x = 100 + -8.33
3.33 x = 91.67
/ 3.33
x= 27.5 mph mini bike I
mini bike II 27.5 + 5 = 32.5 mph
m.ananth@hotmail.ca